Assuming that the covalent radius of hydrogen is 28 pm when it is bonded to other
atoms, and using Table 3.10, calculate the bond lengths of the bonds in
CH2=CH-CH2-NH2​
1
Expert's answer
2020-04-13T02:34:37-0400

Since there is no table for task (3.10), we will use only two data that should probably be there - the bond length CH in methane and NH in ammonia

The bond in methane is the sum of the covalent radii of carbon and hydrogen, the bond in ammonia is the sum of the covalent radii of carbon and nitrogen.

The double bond is slightly smaller (we can assume that the orbitals of sp 2 and sp 3 are proportional to the S contribution, i.e. 33% and 25%, respectively - which means 3/4)


lengths bonds in CH4=110 pm

lengths bonds in NH3=100 pm


covalent radius H = 28 pm

rH+rC=bong CH

rC=110-28=82 pm

rN=100-28=72 pm

bond C-N=rC+rN=154 pm*


bong C=C=(rC+rC)*3/4=120 pm


bond C-C=rC+rC=164 pm

This lengths bond Perhaps they are not extremely accurate, but they allow you to predict and calculate, and, moreover, are consistent with tabular values, spectrophotometry and not only.